OpenStudy (anonymous):

One number is 8 more than a second number. 4 times the first is 14 less than 2 times the 2nd. Find the numbers

sammixboo (sammixboo):

OK so, First (one) number = f Second Number = s So f = 8 + s And 4f = 14 - 2s I believe that is what is it stating. Do you think?